# Toward Efficient Spiking Transformers: Synapse Pruning Meets Synergistic Learning-Based Compensation

### v2

@@ Acknowledgments
Table S5: Configuration Parameters for Experiments
Appendix S3 Additional Results and Analysis
   Pruning sparsity    Sparse weight matrix    Low-rank weight matrix
   Random [%]    $\mathrm{L_{1}P}$ [%]    Random [%]    DSP [%]
   $p=10\%$    93.97    94.72    82.50    93.33
   $p=20\%$    90.74    94.67    77.63    92.26
   $p=30\%$    84.36    94.53    69.16    89.16
   $p=40\%$    71.22    94.56    62.26    81.33
   $p=50\%$    52.68    94.03    47.00    61.76
   $p=60\%$    35.50    92.92    39.77    36.81
   $p=70\%$    31.29    87.72    38.81    28.23
   $p=80\%$    29.03    67.93    23.15    19.02
   $p=90\%$    18.77    34.89    17.42    14.17
   $p=95\%$    13.00    31.74    22.52    18.61
   $p=99\%$    10.49    16.01    17.83    18.88
Table S6: Comparison of models pruned with proposed methods and random pruning.
Figure S9: Visualization of the original and sparse weight matrices. Each pixel represents a value of weight elements and heavy color is correlated with a more higher absolute value.
Figure S10: Visualization of the original and low-rank weight matrices. Each pixel represents a value of weight elements and heavy color is correlated with a more higher absolute value.
Effectiveness of Pruning Methods
To enable efficient model pruning, accurately evaluating the importance of individual components within the model is essential. In the proposed $\mathbf{L_{1}P}$ method, which is designed for constructing sparse weight matrices, the $\mathbf{L_{1}}$ norm is employed to quantify the significance of each matrix element. Elements are then pruned based on their ranking derived from the $\mathbf{L_{1}}$ importance scores. As illustrated in Fig. S9, under a pruning sparsity of $p=90\%$, only a small subset of elements with the highest $\mathbf{L_{1}}$ importance are retained in the sparse weight matrix. To validate the effectiveness of this selection criterion, we conducted a comparative experiment in which elements are pruned randomly (Tab. S6). The results show that models pruned using the random strategy exhibit significantly lower classification accuracy compared to those pruned using the $\mathbf{L_{1}P}$ method. Notably, in the random pruning scenario, a sharp performance drop is observed at a pruning sparsity of $p=30\%$, whereas the same level of degradation occurs at $p=70\%$ with the $\mathbf{L_{1}P}$ approach.
Similarly, in the DSP method, pruning is performed by eliminating dimensions with low significance scores computed using the DVA metric, resulting in a low-rank approximation of the original weight matrix (see Fig. S10). To evaluate the effectiveness of this strategy, we also performed a comparative analysis by randomly pruning dimensions. The model performance with random pruning degrades rapidly, with a noticeable accuracy drop occurring before $p=10\%$, substantially earlier than the degradation point at $p=30\%$ observed with the DSP method. Moreover, across pruning sparsities of $p\leq 60\%$, models pruned with the DSP strategy consistently outperform those with random pruning. Although the post-pruned accuracy in random pruning is slightly higher than that in DSP, this phenomenon is reasonable given the extremely low accuracy.
In summary, the experimental results validate the effectiveness of the proposed $\mathbf{L_{1}P}$ and DSP methods, which offer principled importance metrics for unstructured and structured pruning strategies, respectively.
Detailed Visualization Method and Further Examples
To gain insights into the decision-making process of our ST-based model, we utilize an adapted version of the Attention Rollout method (Abnar and Zuidema 2020). This technique enables us to aggregate attention across all layers and visualize the regions of the input image that significantly influence the model’s predictions. Specifically, for each layer, we compute the fused attention matrix by averaging across the attention heads (Notably, all attention matrices are selected at the final time step). We then incorporate residual connections by adding the identity matrix and normalize each row to ensure it sums to one. The rollout matrix is obtained by recursively multiplying these processed attention matrices from the first to the last layer. Given that our model employs global average pooling for classification instead of a dedicated class token, we derive the importance of each patch by averaging the rollout matrix across all tokens. To emphasize the most salient regions, we apply a thresholding mechanism that sets the smallest attention values to zero based on a predefined discard ratio (set as $0.85$ in our work). Finally, the resulting mask is upsampled to the original image dimensions and overlaid on the input image to produce an intuitive visualization. The entire pipeline is shown in s shown in Algorithm. S1. We also provide more examples in Fig. S11.
Figure S11: Further examples of attention maps as in Figure 6.
Algorithm S1 Attention Rollout Visualization
@@
Supplementary: The Evolutionary Process of IP Parameters
The evolution of IP parameters for the $Q$, $K$, and $V$ embeddings is illustrated in Fig. S12. Overall, the magnitude of IP parameter adaptation exhibits a positive correlation with pruning sparsity across all embeddings. Notably, this evolution is more pronounced on the CIFAR10-DVS dataset. These observations indicate that the influence of intrinsic plasticity is consistently reflected across the ST blocks.
Qualitative Results on the ADE20K Dataset
@@
As illustrated in Fig. S13, the semantic segmentation performance of the lightweight model exhibits a slight degradation compared to the original model. Nonetheless, considering that the lightweight model utilizes only half the number of parameters, its performance remains competitive and acceptable for practical applications.
Appendix S4 Limitations
Our compression strategy offers flexibility in obtaining lightweight models by allowing manual configuration of the predefined pruning sparsity. However, the same pruning sparsity is applied uniformly across all ST blocks, potentially overlooking the varying contributions of different blocks to the overall model performance. As part of future work, a comprehensive evaluation metric should be developed to enable dynamic adjustment of pruning levels across ST blocks based on their relative importance. Hybrid models that integrate diverse architectural components, such as convolutional blocks, MLP blocks, and transformer blocks, have demonstrated increasing potential in addressing complex tasks. Accordingly, extending our method to support a broader range of architectures may further enhance its applicability and impact across various model designs and domains.
@@
Figure S12: Evolution of IP parameters of the Q, K and V embeddings, respectively. ”Start” indicates the initial parameter values.
Figure S13: Qualitative semantic segmentation results on the ADE20K dataset. The top row depicts ground truth masks, while the middle and bottom rows show the predicted masks produced by the original and lightweight models, respectively.

### v4

@@ III-C2 Response Re-alignment via Synergistic Plasticity
TABLE II: Classification results on the ImageNet-100 dataset. The symbol denotes the baseline model. The terms p and CR indicate the pruning sparsity and compression ratio, respectively. The notation ‘/’ separates values measured on the entire model and on the ST blocks.
Method Architecture p [%] CR [%] Param [M] Accuracy [%]
LOCALZO+TET [mukhoty2023direct] SEWResNet34 - - 63.47 78.58
IM-SNN [hasssan2024snn] Resnet34 - - 21.27 74.42
IMP+TET [shen2024rethinking] SEW-ResNet18 - - 63.47 78.70
EfficientLIF-Net [10.3389/fnins.2023.1230002] VGG16 - - 23.52 73.22
Spikformer [zhou2023spikformer] Spikformer-8-512-2048 - - 29.24/25.17 79.36
TABLE III: Classification results on the CIFAR10 and CIFAR10-DVS datasets. The symbol denotes the baseline model. The terms p and CR indicate the pruning sparsity and compression ratio, respectively. The notation ‘/’ separates values measured on the entire model and on the ST blocks.
Dataset Method Architecture p [%] CR [%] Param [M] Accuracy [%]
CIFAR10 STBP-tdBN [zheng2021going] ResNet-19 - - 11.17 92.92
STP [ma2025spiking] ResNet18 - - 63.47 94.86
STL-SNN [sun2023synapse] ConvFC - - 11.37 92.42
Spikformer [zhou2023spikformer] Spikformer-4-384-1536 - - 9.32/7.08 94.76
Spikingformer [spikingformer] Spikformer-4-384-1536 - - 9.32/7.08 95.81
CML [zhou2023enhancing] Spikformer-4-384-1536 - - 9.32/7.08 96.04
$\mathrm{L_{1}P}$+sLIF (ours) Spikformer-4-384-1536 20 18.03/23.73 7.64/5.40 94.72(-0.04)
Spikformer-4-384-1536 40 32.51/42.80 6.29/4.05 94.78(+0.02)
Spikformer-4-384-1536 60 47.00/61.86 4.94/2.70 94.61(-0.15)
Spikformer-4-384-1536 80 60.62/79.94 3.67/1.42 93.94(-0.82)
Spikformer-4-384-1536 90 68.24/89.97 2.96/0.71 92.32(-2.44)
DSP+sLIF (ours) Spikformer-4-312-1228 20 17.60/23.31 7.68/5.43 94.60(-0.16)
Spikformer-4-240-921 40 31.76/41.95 6.36/4.11 94.36(-0.40)
Spikformer-4-156-614 60 46.67/61.58 4.97/2.72 94.03(-0.73)
Spikformer-4-84-307 80 60.19/79.38 3.71/1.46 93.14(-1.62)
Spikformer-4-48-153 90 67.60/89.12 3.02/0.77 92.23(-2.53)
CIFAR10-DVS TET [deng2022temporal] VGGSNN - - 9.54 77.33
STP [ma2025spiking] VGG11 - - 113.00 78.50
STL-SNN [sun2023synapse] ConvFC - - 1.53 77.30
Spikformer [zhou2023spikformer] Spikformer-2-256-1024 - - 2.59/1.58 79.40
Spikingformer [spikingformer] Spikformer-2-256-1024 - - 2.59/1.58 81.30
CML [zhou2023enhancing] Spikformer-2-256-1024 - - 2.57/1.58 79.20
$\mathrm{L_{1}P}$+sLIF (ours) Spikformer-2-256-1024 20 14.67/24.05 2.21/1.20 80.00(+0.60)
Spikformer-2-256-1024 40 26.25/43.04 1.91/0.90 80.10(+0.70)
Spikformer-2-256-1024 60 37.84/62.03 1.61/0.60 80.40(+1.00)
Spikformer-2-256-1024 80 48.65/79.75 1.33/0.32 78.00(-1.40)
Spikformer-2-256-1024 90 54.83/25.95 1.17/0.16 76.30(-3.10)
DSP+sLIF (ours) Spikformer-2-192-819 20 15.83/43.67 2.18/1.17 78.90(-0.50)
Spikformer-2-144-614 40 26.64/62.66 1.90/0.89 78.90(-0.50)
Spikformer-2-96-409 60 38.22/80.38 1.60/0.59 78.90(-0.50)
Spikformer-2-48-204 80 49.03/80.38 1.32/0.31 77.20(-2.20)
Spikformer-2-16-102 90 55.60/91.14 1.15/0.14 77.30(-2.10)
IV Experiments and Results
IV-A Experimental Settings
To evaluate the effectiveness of our method, experiments are conducted on both static and neuromorphic image datasets, ImageNet [5206848], CIFAR-10 [krizhevsky2009learning], and CIFAR10-DVS [li2017cifar10]. To further validate the generalization capability in downstream tasks, semantic segmentation is performed on the ADE20K dataset [zhou2017scene]. For a fair comparison, our method are directly applied to state-of-the-art (SOTA) pre-trained models implemented using the PyTorch and SpikingJelly frameworks [zhou2023spikformer, yao2024spikedriven, doi:10.1126/sciadv.adi1480]. All experiments are executed on 4 NVIDIA A800 GPUs.
IV-A1 Datasets
ImageNet [5206848] is a large-scale static image dataset widely used for object recognition tasks, comprising approximately 1.28 million training images and 50,000 validation images across 1,000 classes. In our experiments, all images are resized to a resolution of $\times224$ pixels for consistency. To evaluate the classification performance of our method on high-resolution static images, we adopt the ImageNet-100 subset [imagenet100pytorch], where the classes are selected following prior work. This subset serves as a benchmark for comparison with state-of-the-art (SOTA) models in conventional computer vision tasks. Additionally, the full ImageNet-1K dataset is employed to assess the scalability of our method under the scaling-law setting.
CIFAR-10 [krizhevsky2009learning] is a static image dataset consisting of 50,000 training images and 10,000 test images, each with a resolution of $\times32$ pixels and evenly distributed across 10 classes. It is used to evaluate the effectiveness of our method on low-resolution static images, offering a lightweight yet challenging benchmark for classification tasks.
CIFAR10-DVS [li2017cifar10] is a neuromorphic dataset derived from the original CIFAR-10, recorded using a dynamic vision sensor (DVS). It comprises event streams from 10,000 samples spanning the same 10 classes as CIFAR-10, where each sample is represented as a sequence of address-event representations (AER) instead of conventional image frames. With temporal resolution on the order of microseconds, this dataset is well-suited for evaluating the performance of our method on event-based, neuromorphic vision tasks. In our experiments, each event stream is temporally averaged and segmented into 16 discrete time steps.
ADE20K [zhou2017scene] is a widely used dataset for semantic segmentation, comprising 20,210 training images and 2,000 validation images annotated with 150 semantic categories. For segmentation tasks, images are typically resized to a resolution of $\times512$ pixels. Owing to its diverse scene compositions and fine-grained annotations, ADE20K serves as a benchmark to assess the generalization capability of our method on downstream tasks, particularly semantic segmentation.
IV-A2 Configuration Details
Tab. I summarizes the key configuration parameters used for each dataset in our experiments. The source code will be released upon acceptance at https://github.com/GuoLab-UESTC/EfficientST.
Figure 5: Performance comparison of the proposed lightweight models (sLIF) and LIF-compensated models (LIF) under varying pruning sparsity on the CIFAR10 and CIFAR10-DVS datasets. The parameter counts (Param) and pruned accuracies (Pruned) of the baseline models are also provided to demonstrate the effectiveness of the proposed method.
IV-B Comparisons with SOTA Methods
IV-B1 Static Datasets Classification
We compare our lightweight models with other SNNs on the ImageNet-100 dataset (Tab. II). The baseline Spikformer-8-512-2048 model consists of eight ST blocks with $d=512$ and $d_{\mathrm{m}}=2048$. After pruning with a sparsity level of $p=90\%$, the parameter count is reduced to below 77M, achieving a compression ratio (CR) exceeding $77\%$, while incurring only a minor accuracy drop ($3.14\%$ for $\mathrm{L_{1}P}$+sLIF and $2.48\%$ for DSP+sLIF). Considering only the ST blocks, the CR exceeds $90\%$, with parameters reduced to below 2.52.5M. When sparsity is further increased to $p=99\%$, the accuracy decreases by about $16\%$. Nevertheless, given that the parameter count of the ST blocks is reduced from 2424M to only 0.250.25M (for $\mathrm{L_{1}P}$+sLIF) or 0.240.24M (for DSP+sLIF), this trade-off remains acceptable for lightweight deployment. Moreover, compared to ResNet- or VGG-based convolutional models with comparable accuracy, our lightweight models demonstrate substantially higher parameter efficiency.
On the CIFAR10 dataset (Tab. III), ST-based models such as Spikformer and CML exhibit superior performance, with CML attaining the highest accuracy of $96.04\%$ when employing the Spikformer-4-384-1536 architecture. Utilizing our proposed $\mathrm{L_{1}P}$+sLIF method, an accuracy of $93.94\%$ is achieved under a sparsity level of $p=80\%$, accompanied by a substantial compression ratio of $60.62\%$ for the entire model and $79.94\%$ for the ST modules, and a significantly reduced parameter count of 3.673.67M and 1.421.42M, respectively. When the sparsity is further increased to $p=90\%$, the model maintains a competitive accuracy of $92.32\%$, while the parameter count is further reduced to 2.962.96M (entire model) and 0.710.71M (ST modules). Similarly, the proposed DSP+sLIF strategy also demonstrates strong performance with enhanced parameter efficiency, achieving $93.14\%$ ($92.23\%$) accuracy with only 3.713.71M (3.023.02M) parameters under $p=80\%$ ($p=90\%$).
IV-B2 Neuromorphic Datasets Classification
On the CIFAR10-DVS dataset, the proposed lightweight models also demonstrate competitive performance, as summarized in Tab. III. The baseline Spikformer model achieves an accuracy of $80.90\%$ with 2.592.59M parameters (1.581.58M in ST modules). By applying the $\mathrm{L_{1}P}$+sLIF method under a sparsity level of $p=80\%$, the model is compressed to 1.331.33M parameters (0.320.32M in ST modules) while maintaining a commendable accuracy of $78.00\%$. Similarly, the DSP+sLIF strategy yields a compact Spikformer-2-48-204 model with 1.321.32M parameters (0.310.31M in ST modules) and an accuracy of $77.20\%$ at the same sparsity level. Performance remains comparable when sparsity is further increased.
Figure 6: Comparison of convergence curves of pruned models with LIF- and sLIF-based compensation on the CIFAR-10 dataset during fine-tuning phase.
Figure 8: Representative examples of attention maps generated by the original (middle) and lightweight (right) models on the ImageNet-100 dataset. Higher attention values are indicated in deep red.
Algorithm 1 Attention Rollout Visualization
@@ IV-B2 Neuromorphic Datasets Classification
IV-C Ablation Analysis
IV-C1 Performance Under Different Levels of Sparsity
To comprehensively evaluate the performance of the proposed lightweight strategies, we perform pruning experiments under different sparsity levels (denoted as ‘sLIF’). For comparison, we also assess the accuracy of the original model after pruning (denoted as ‘Pruned’). Additionally, to evaluate the contribution of synergistic learning, we conduct pruning experiments while retaining the original LIF neurons (denoted as ‘LIF’). As illustrated in Fig. 5, model accuracy generally declines after pruning, with greater sparsity levels resulting in more pronounced performance degradation. Compared to the unstructured pruning strategy ($\mathrm{L_{1}P}$), the structured pruning approach (DSP) causes more substantial performance deterioration, which is consistent with prior findings. Nonetheless, fine-tuning is consistently an efficient strategy to recover performance. Notably, models fine-tuned with synergistic learning exhibit superior performance recovery, with the benefits of synergistic learning becoming increasingly evident at higher sparsity levels.
In addition to superior compensation performance, the convergence speed during the fine-tuning process is a critical factor in evaluating the effectiveness of a compensation strategy. We present the accuracy curves for the ‘sLIF’ and ‘LIF’ models under pruning sparsity of $p=99\%$ and $p=95\%$ on the CIFAR10 dataset, as shown in Fig. 6. In our experiments, we configured the fine-tuning process to run for 50 epochs. Across models pruned with various strategies, the sLIF approach consistently enables faster convergence during fine-tuning. Notably, this advantage is particularly pronounced for models pruned using the DSP method. Specifically, models fine-tuned with sLIF achieve convergence in approximately 20 epochs, whereas the ablation group employing LIF neurons requires around 40 epochs to reach convergence.
IV-C2 Impact on Model Inference Performance
Inference latency and energy consumption represent two fundamental metrics for assessing the inference performance of models. Since structured pruning facilitates improved hardware accessibility, we evaluate the average batch inference runtime and theoretical power consumption of the DSP+LIF and DSP+sLIF models under pruning sparsity levels of $p=50\%$, $90\%$, and $99\%$, respectively. The relative ratios of the lightweight models compared to their corresponding baseline models are presented in Fig. 7.
SNN models predominantly utilize sparse accumulate operations as their primary computational units, resulting in significantly lower power consumption compared to traditional artificial neural networks. Additionally, the overall energy consumption of SNNs is influenced by both the model architecture and the neuronal firing rate. In this study, we estimate theoretical energy consumption following the method used in prior work [sun2024reliable]. As shown in Fig. 7, the relative energy consumption of pruned models, compared to the baseline, shows a positive correlation with the number of model parameters. Furthermore, sLIF models demonstrate slightly higher energy consumption than their LIF counterparts.
TABLE IV: Performance on the large-scale ImageNet-1k dataset.
Sparsity baseline $p=30\%$ $p=50\%$ $p=90\%$
Param [M] 29.71 22.05 17.12 7.00
Acc [%] 72.86 67.70 65.02 56.00
IV-C3 Visualization of Attention Maps
To gain insight into the decision-making process of our ST-based model, we utilize an adapted version of the Attention Rollout method [Abnar2020QuantifyingAF]. This technique enables us to aggregate attention across all layers and visualize the regions of the input image that significantly influence the predictions of the model. Specifically, for each layer, we compute the fused attention matrix by averaging across the attention heads (Notably, all attention matrices are selected at the final time step). We then incorporate residual connections by adding the identity matrix and normalize each row to ensure it sums to one. The rollout matrix is obtained by recursively multiplying these processed attention matrices from the first to the last layer. Given that our model employs global average pooling for classification instead of a dedicated class token, we derive the importance of each patch by averaging the rollout matrix across all tokens. To emphasize the most salient regions, we apply a thresholding mechanism that sets the smallest attention values to zero based on a predefined discard ratio (set to $0.85$ in our work). Finally, the resulting mask is upsampled to the original image dimensions and overlaid on the input image to produce an intuitive visualization. The entire pipeline is shown in Algorithm 1.
In the DSP+sLIF models, the dimensions of the $Q$, $K$, and $V$ representations are reduced to extremely low levels. For instance, pruning the original ‘Spikformer-8-512-2048’ architecture (Tab. II) with a pruning sparsity of $p=90\%$ results in a lightweight ‘Spikformer-8-48-204’ model, where only 48 dimensions are used to represent the patch embeddings. To assess the representational capacity of the lightweight model, we compare the attention maps at the final time step between the baseline (‘Spikformer-8-512-2048’) and pruned models (‘Spikformer-8-48-204’) on the ImageNet-100 dataset (shown in Fig. 8). Despite the reduced dimensionality, the lightweight model still effectively captures image regions relevant to classification tasks. These results demonstrate that our compensation strategy is an efficient and effective approach for compressing ST-based models.
TABLE V: Semantic segmentation performance on the ADE20K dataset.
Method ST p [%] Param [M] mIoU [%] mAcc [%]
MetaFormer [yu2022metaformer] $\times$ - 15.50 32.90 -
DeeplabV3 [zhang2022resnest] $\times$ - 68.10 42.10 -
SDTv2 [yao2024spikedriven] $\checkmark$ - 9.42 30.14 42.22
DSP+sLIF (ours) $\checkmark$ 50 6.52 29.71 41.65
DSP+sLIF (ours) $\checkmark$ 75 5.36 27.60 38.51
DSP+sLIF (ours) $\checkmark$ 90 4.70 26.69 37.73
Figure 9: Qualitative semantic segmentation results on the ADE20K dataset.
IV-D Performance on the Large-Scale Dataset
We conduct experiments with the DSP+sLIF method on the more challenging ImageNet-1K dataset (Tab. IV). The model architecture and experimental settings are aligned with those used for ImageNet-100. Compared to ImageNet-100, a more significant performance drop is observed at high sparsity levels (e.g., $16.86\%$ on ImageNet-1K vs. $2.48\%$ on ImageNet-100 at $p=90\%$). This degradation is theoretically reasonable, as larger-scale datasets require larger model size according to scaling laws [dosovitskiy2021an]. Nevertheless, the DSP+sLIF method remains effective, achieving competitive performance under extreme compression even on large-scale datasets.
IV-E Performance on the Downstream Tasks
To further demonstrate the effectiveness of the proposed DSP+sLIF method on complex downstream tasks, we apply it to semantic segmentation. The ST-based SDTv2 models [yao2024spikedriven] are used as the baseline, and the DSP+sLIF strategy is employed to compress the original models with pruning sparsity levels of $p=50\%$, $p=75\%$, and $p=90\%$. As shown in Tab. V, the resulting lightweight ST-based semantic segmentation models achieve $26.69\%$ mIoU and $37.73\%$ mAcc performance while requiring only 4.704.70M parameters. Moreover, appropriately enlarging the size of the lightweight models further mitigates the performance loss. When the baseline model is compressed with a sparsity level of $p=50\%$, the mIoU (mAcc) decreases only slightly by $0.43\%$ ($0.57\%$).
@@ IV-E Performance on the Downstream Tasks
As illustrated in Fig. 9, the semantic segmentation performance of the lightweight model ($p=90\%$) exhibits a slight degradation compared to the original model. Nonetheless, considering that the lightweight model utilizes only half the number of parameters, its performance remains competitive and acceptable for practical applications.
Figure 10: Visualization of the original, sparse weight matrices and low-rank weight matrices. Each pixel represents a value of weight elements and heavy color is correlated with a more higher absolute value.
TABLE VI: Comparison of models pruned with proposed methods and random pruning.
p [%] Sparse weight matrix Low-rank weight matrix
Random [%] $\mathrm{L_{1}P}$ [%] Random [%] DSP [%]
$10$ 93.97 94.72 82.50 93.33
2020 90.74 94.67 77.63 92.26
$30$ 84.36 94.53 69.16 89.16
$40$ 71.22 94.56 62.26 81.33
$50$ 52.68 94.03 47.00 61.76
$60$ 35.50 92.92 39.77 36.81
$70$ 31.29 87.72 38.81 28.23
$80$ 29.03 67.93 23.15 19.02
$90$ 18.77 34.89 17.42 14.17
$95$ 13.00 31.74 22.52 18.61
$99$ 10.49 16.01 17.83 18.88
@@
Figure 11: Evolution of IP parameters of the Q, K and V embeddings under different pruning sparsity. ‘Start’ indicates the initial parameter values.
Figure 12: (a) Firing rate comparison of pruned models with LIF- and sLIF-based compensation in the ST modules. The black dotted line indicates the firing rate of the baseline model. (b) Probability density distribution of the input current in $x_{\mathrm{attn}}$ spiking layers.
IV-F Effectiveness of Pruning Methods
To enable efficient model pruning, accurately evaluating the importance of individual components within the model is essential. In the proposed L1P method, which is designed for constructing sparse weight matrices, the L1P norm is employed to quantify the significance of each matrix element. Elements are then pruned based on their ranking derived from the L1P importance scores. As illustrated in Fig. 10, under a pruning sparsity of $p=90\%$, only a small subset of elements with the highest L1 importance is retained in the sparse weight matrix. To validate the effectiveness of this selection criterion, we conducted a comparative experiment in which elements were pruned randomly (Tab. VI). The results show that models pruned using the random strategy exhibit significantly lower classification accuracy compared to those pruned using the L1P method. Notably, in the random pruning scenario, a sharp performance drop is observed at a pruning sparsity of $p=30\%$, whereas the same level of degradation occurs at $p=70\%$ with the L1P approach.
Similarly, in the DSP method, pruning is performed by eliminating dimensions with low significance scores computed using the DVA metric, resulting in a low-rank approximation of the original weight matrix (see Fig. 10). To evaluate the effectiveness of this strategy, we also performed a comparative analysis by randomly pruning dimensions. The model performance with random pruning degrades rapidly, with a noticeable accuracy drop occurring before $p=10\%$, substantially earlier than the degradation point at $p=30\%$ observed with the DSP method. Moreover, across pruning sparsities of $p\leq 60\%$, models pruned with the DSP strategy consistently outperform those with random pruning. Although the post-pruned accuracy in random pruning is slightly higher than that in DSP, this phenomenon is reasonable given the extremely low accuracy.
In summary, the experimental results validate the effectiveness of the proposed L1P and DSP methods, which offer principled importance metrics for unstructured and structured pruning strategies, respectively.
IV-G Mechanism of Synergistic Learning-Based Compensation
To evaluate the contribution of synergistic learning to model performance, we visualize the evolution of the IP parameters, $u_{\mathrm{th}}$ and $\tau$, associated with the $Q$, $K$ and $V$ embeddings in the first ST block on the CIFAR10 and CIFAR10-DVS datasets. As illustrated in Fig. 11, the magnitude of IP parameter adaptation is positively correlated with the sparsity level. Moreover, the evolution is more pronounced on the more challenging CIFAR10-DVS dataset. This phenomenon is consistent with the theoretical analysis from the perspective of neural dynamics, indicating that IP parameters in synergistic learning play a significant role in enhancing model performance under increased task difficulty.
TABLE VII: Accuracy comparison of lightweight models compensated with various IP mechanisms.
Neuron $\mathrm{L_{1}P}$ DSP
$5\% p=99p$$9\%$ $5\% p=99p$$9\%$
LIF 81.80% 64.60% 83.74% 77.93%
PLIF [Fang_2021_ICCV] 87.63% 71.96% 89.76% 86.03%
STL [sun2023synapse] 87.75% 72.30% 89.79% 86.00%
sLIF 88.86% 74.93% 90.64% 87.94%
In SNN models, maintaining an appropriate firing rate is essential for preserving information representation and ensuring stable spike propagation. This requirement is particularly critical in lightweight models, where a moderate increase in firing rate enables a larger population of neurons to participate in encoding, thereby reducing redundancy and mitigating information loss. As illustrated in Fig. 12(a), the firing rates of pruned models with LIF-based compensation progressively decrease as pruning sparsity increases. In contrast, models employing sLIF-based compensation consistently restore firing rates to baseline levels. Notably, at an extreme pruning sparsity of $p=99\%$, the firing rate slightly exceeds the baseline, aligning with expectations. Although the elevated firing rate may incur marginally higher power consumption, this overhead remains acceptable given the performance gains achieved by the post-compensated models. Furthermore, we analyze the probability density distribution of input currents in the $x_{\mathrm{attn}}$ spiking layers, generated by the scaled dot-product attention mechanism (Fig. 12(b)). At higher pruning sparsity levels ($p=70\%$ and $p=99\%$), clear differences emerge between LIF- and sLIF-based compensation. The synergistic learning mechanism produces wider distributions with larger expectations, underscoring its role in post-pruning information compensation.
Additionally, we conduct experiments to compensate pruned models using different intrinsic plasticity mechanisms. Prior studies have enhanced neuronal expressiveness by optimizing intrinsic properties, such as the membrane time constant $\tau$ (PLIF) [Fang_2021_ICCV] or the firing threshold $u_{\text{th}}$ (STL) [sun2023synapse]. However, compensation based on synergistic learning achieves more superior performance, as reported in Tab. VII. As task difficulty increases, constructing more detailed neuron models becomes an effective means of improving the performance of SNNs [CAI2026108417, pnas2513319122, liu2024dendritic]. Compared with these approaches, synergistic learning provides a simple yet efficient alternative, yielding improved training performance without introducing additional computational overhead during the inference stage.
V Discussion and Conclusion
From a biological perspective, synaptic plasticity regulates information transmission by adjusting the strength of inter-neuronal connections, whereas intrinsic plasticity modifies neuronal properties such as membrane time constants and firing thresholds to control excitability. In the brain, these mechanisms operate synergistically to maintain network stability and robustness. When pruning eliminates a substantial portion of synaptic connections, relying solely on synaptic adjustment often results in reduced firing rates and diminished representational capacity. By jointly adapting intrinsic parameters, synergistic learning preserves population firing rates and increases neuronal heterogeneity, thereby enhancing the diversity of temporal and spatial representations. This enables compressed models to encode information effectively across multiple timescales and feature dimensions, even under high sparsity. Such a mechanism parallels the compensatory strategies observed in biological neural systems under damage or resource constraints, explaining why synergistic learning provides more comprehensive and robust information compensation in pruned models. Consistently, our experimental results demonstrate that synergistic learning facilitates superior performance recovery in heavily pruned networks.
The limitations and future work of the proposed compression strategy are discussed below. Our compression strategy offers flexibility in obtaining lightweight models by allowing manual configuration of the predefined pruning sparsity. However, the same pruning sparsity is applied uniformly across all ST blocks, potentially overlooking the varying contributions of different blocks to the overall model performance. As part of future work, a comprehensive evaluation metric should be developed to enable dynamic adjustment of pruning levels across ST blocks based on their relative importance. Hybrid models that integrate diverse architectural components, such as convolutional blocks, MLP blocks, and transformer blocks, have demonstrated increasing potential in addressing complex tasks. Accordingly, extending our method to support a broader range of architectures may further enhance its applicability and impact across various model designs and domains.